OpenStudy (anonymous):

4x + 3y = 9... using y =mx +B

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

What do you want to do?

OpenStudy (anonymous):

m=-4/3, want the slope?

OpenStudy (gw2011):

If you want it written in the form of: y=mx+b, then it becomes: 3y=-4x+9 y=-4/3x+3